OpenStudy (anonymous):

Help: I need help with Linear Equations: 5(x-3)+4x-1=2x+3(x-2)

OpenStudy (anonymous):

Find x?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Distribute 5, also 3 into the respective bracket, can you?

OpenStudy (anonymous):

Ill show you what i have and can you tell me where i went wrong?

OpenStudy (anonymous):

This is the original question: 5(x-3)+4x-1=2x+3(x-2) This is what i got next: 5x-15+4x-1=2x+3x-6 Then i got this: 9x-16=5x-6 Then i got this: 9x+5x=-22 Is this right?

OpenStudy (anonymous):

That last one is wrong, you need to flip the signs when you switch sides...

OpenStudy (anonymous):

So what would i flip? -16 to 16 ?

OpenStudy (anonymous):

What I mean is if you bring something from one side of the equals to the other then you need to change the sign of that something.....

OpenStudy (anonymous):

Or, the way they teach you to do it is like: 9x-16=5x-6 now subtract 5x from both sides -> 9x-16 -5x = 5x - 6 -5x 4x -16 v= -6 now add 16 to both sides etc etc

OpenStudy (anonymous):

Comes to the same thing...

OpenStudy (anonymous):

so would your answer be x=2.5? cause you divide the 10 by 4?

OpenStudy (anonymous):

4x -16 v= -6 4x = 10 Now divide both sides by 4 x = 5/2 (or 2.5 if you like)

OpenStudy (anonymous):

Okay thank you so much!